# A Transformer approach for Electricity Price Forecasting

### v1

@@ Abstract
This paper presents a novel approach to electricity price forecasting (EPF) using a pure Transformer model. As opposed to other alternatives, no other recurrent network is used in combination to the attention mechanism. Hence, showing that the attention layer is enough for capturing the temporal patterns. The paper also provides fair comparison of the models using the open-source EPF toolbox and provide the code to enhance reproducibility and transparency in EPF research. The results show that the Transformer model outperforms traditional methods, offering a promising solution for reliable and sustainable power system operation [^1] All code to replicate our findings will be available here: https://github.com/osllogon/positive_active_saliency_maps.
@@ II Proposed Transformer model for Electricity Price Forecasting
The main model architecture is illustrated in  1(a), which consists of two main data flows, one for each input. Then, the results of each flow are concatenated and served as an input to a multivariate Multilayer Perceptron for predicting day ahead prices. The architecture is described ahead in detail.
@@ IV Results
MAE
RMSE
sMAPE
Nord Pool
2.89
5.31
0.08
PJM
4.43
7.24
0.17
EPEX-BE
9.38
20.97
0.23
EPEX-FR
7.24
15.2
0.22
EPEX-DE
8.19
13.19
0.3
@@
MAE
RMSE
sMAPE
Nord Pool
3.4
5.27
0.09
PJM
7.17
10.24
0.27
EPEX-BE
9.38
18.49
0.24
EPEX-FR
4.74
12.11
0.14
EPEX-DE
4.34
7.63
0.17
@@
MAE
RMSE
sMAPE
Nord Pool
3.78
5.73
0.09
PJM
11.98
17.92
0.35
EPEX-BE
10.87
20.68
0.28
EPEX-FR
4.58
11.99
0.13
EPEX-DE
4.38
8.07
0.17
@@
Nord Pool
2.33
4.08
0.07
0
0
0
PJM
3.67
5.85
0.14
0
0
0
EPEX-BE
6.54
16.68
0.15
0
0
0
EPEX-FR
4.91
12.67
0.14
0.97
1
0
EXPEX-DE
4.03
6.99
0.17
0
0
0

### v2

@@ Abstract
This paper presents a novel approach to electricity price forecasting (EPF) using a pure Transformer model. As opposed to other alternatives, no other recurrent network is used in combination to the attention mechanism. Hence, showing that the attention layer is enough for capturing the temporal patterns. The paper also provides fair comparison of the models using the open-source EPF toolbox and provide the code to enhance reproducibility and transparency in EPF research. The results show that the Transformer model outperforms traditional methods, offering a promising solution for reliable and sustainable power system operation [^1] All code to replicate our findings will be available here: https://github.com/osllogon/epf-transformers.
@@ II Proposed Transformer model for Electricity Price Forecasting
The main model architecture is illustrated in  2(a), which consists of two main data flows, one for each input. Then, the results of each flow are concatenated and served as an input to a multivariate Multilayer Perceptron for predicting day ahead prices. The architecture is described ahead in detail.
@@ IV Results
MAE RMSE sMAPE
Nord Pool 2.89 5.31 0.08
PJM 4.43 7.24 0.17
EPEX-BE 9.38 20.97 0.23
EPEX-FR 7.24 15.2 0.22
EPEX-DE 8.19 13.19 0.3
@@
MAE RMSE sMAPE
Nord Pool 3.4 5.27 0.09
PJM 7.17 10.24 0.27
EPEX-BE 9.38 18.49 0.24
EPEX-FR 4.74 12.11 0.14
EPEX-DE 4.34 7.63 0.17
@@
MAE RMSE sMAPE
Nord Pool 3.78 5.73 0.09
PJM 11.98 17.92 0.35
EPEX-BE 10.87 20.68 0.28
EPEX-FR 4.58 11.99 0.13
EPEX-DE 4.38 8.07 0.17
@@
Nord Pool 2.33 4.08 0.07 0 0 0
PJM 3.67 5.85 0.14 0 0 0
EPEX-BE 6.54 16.68 0.15 0 0 0
EPEX-FR 4.91 12.67 0.14 0.97 1 0
EXPEX-DE 4.03 6.99 0.17 0 0 0